# Human-in-the-Loop Optimization with Model-Informed Priors

### v1

@@ 3. Human-in-the-Loop Optimization with Model-Informed Priors
Our first contribution is the introduction of a conceptual framework, which we term HOMI. We distinguish HOMI from our proposed method (detailed in the next section): While NAF+ represents one possible realization, the HOMI framework is intended to be general and extensible, enabling future work to explore alternative implementations. Below, we outline the core components of the framework.
@@ 3.1. Core Elements of HOMI:
The HOMI framework consists of three key components: meta-Bayesian optimization (meta-BO), user models, and end-users. Meta-BO refers to a computational optimizer trained across a distribution of synthetic users for the same interaction (or application). During the optimizer training phase, meta-BO interacts with synthetic users generated from selected, relevant user models. Since these interactions are simulated, the optimizer can perform an unlimited number of evaluations to learn generalized adaptation strategies. During the deployment phase, meta-BO engages with real end-users. Its goal is to quickly infer user-specific preferences or performance characteristics, and identify optimal designs with minimal interactions. User models are parameterized functions (e.g., Fitts’ law) that simulate user behavior and performance. By sampling different model parameter settings, we can generate a diverse population of synthetic users, each representing different motor abilities, interaction speeds, or usage contexts. End-users are real human participants interacting with the system during deployment. The meta-BO observes real-time signals, such as performance metrics, interaction preferences, or behavioral data—and uses them to adaptively select new design candidates.
@@ 3.2. General Steps in HOMI:
Figure 1 illustrates the four main steps in the HOMI framework. Note that the first three steps comprise the offline training phase, while only the fourth step occurs during real-time deployment.
Step 1: Model selection. HOMI framework requires synthetic users with varying performances to train the optimizer. User models are the tool for this. In HCI, many models are proposed to capture and predict user performances for different interactions and contexts (Purificato et al., 2024); by varying the parameter settings within these models, different user performance can be simulated. Thus, the first step of our framework is to identify the models relevant to the target task; that is, a model that can consider the variation in design and predict different user performances accordingly. For example, for typing and target selection, Fitt’s law is the established model for predicting the movement time. These user models often require parameter tuning to accurately predict performance in a given setting or context. This step may involve fitting model parameters to empirical data or adopting values reported in prior literature, yielding relevant parameter distributions. This is optional: if reliable parameter estimates already exist, they can be directly reused. For instance, touchscreen touch accuracy has been modeled using a 2D Gaussian distribution with well-documented parameter values (Bi and Zhai, 2013), which can serve as a ready basis for generating synthetic users. There may be cases that there is not a relevant model yet for a novel interaction technique. In such cases, introducing novel user models will be needed.
Step 2: Synthetic user generation: After acquiring reliable model parameter estimates and ranges, we can then sample possible model parameters from the parameter distributions. This leads to a diverse set of synthetic users, each simulating distinct user characteristics.
Step 3: Meta-BO training with synthetic interactions: Before deploying to the end-users, meta-BO interacts with these synthetic users to learn how to efficiently identify optimal designs across different user profiles. Since synthetic evaluations are inexpensive, compared to human evaluations, this phase allows for large-scale training. The diversity of synthetic users further ensures the optimizer does not overfit to specific traits, but rather learns transferrable strategies for personalization.
@@ 4.1. Preliminaries
To better introduce our proposed method, we briefly review three foundations: standard acquisition functions and their limitations, reinforcement learning, and the Neural Acquisition Function (NAF). Our NAF+ builds directly on these concepts and implementations.
1. Acquisition functions in BO
As reviewed in Related Work, BO has two core components: a surrogate model that approximates the target objective function (typically, GP) and an acquisition function that guides the search for the next design candidate to evaluate. The primary goal of BO is to efficiently identify the optimal point (i.e., design) with a minimal number of evaluations. The acquisition function leverages information from the surrogate model (specifically, the predicted mean and predicted variance for each candidate design) to compute an acquisition value. This value represents the potential of evaluating that candidate in the next iteration. The acquisition function is computed across the design space, and the candidate with the highest acquisition value is selected for evaluation. However, a shared limitation of the conventional acquisition functions is that they are not task-specific: their heuristics are fixed and do not adapt based on prior experience or the characteristics of a given domain. This limits their efficiency, especially in human-involved tasks where evaluation cost is high and adaptation must be fast. This motivates our transition toward learned acquisition functions, where the search strategy itself is learned from prior data–collected from other similar users or, in our case, synthetic users–tailored to the structure of the problem and capable of generalizing across users or tasks.
@@ 2. Deep reinforcement learning
The core idea of Neural Acquisition Function (NAF) is to learn a task-specific acquisition function using a neural network. Instead of relying on hand-designed heuristics like Expected Improvement or UCB, NAF aims to encode an adaptive search strategy that is tailored to a particular class of optimization tasks. A key challenge in learning such an acquisition function is the absence of ground truth labels — i.e., there is no “correct” acquisition value to supervise the training. The predictions are inherently uncertain, noisy, and task-dependent. To overcome this, NAF is trained using reinforcement learning (RL). Specifically, the neural network generates acquisition scores across the design space. The design candidate with the highest score is selected for evaluation. The outcome of this evaluation — i.e., the observed objective function value — is then converted into a reward signal: promising evaluation results yield positive rewards, while poor results generate negative feedback. Over time, the model learns to associate high acquisition values with promising candidates from current status, refining its search policy. Viewed from an RL perspective, meta-BO acts as an RL agent, and the neural acquisition function (NAF) serves as its policy model, determining which action (i.e., which design candidate) the agent should take in each iteration given the GP’s status. This formulation enables task-level learning, where the acquisition policy improves through repeated interaction with synthetic training tasks. Below, we briefly review the concept of RL and the corresponding method for training NAF.
@@
RL is a machine learning paradigm that trains an agent to develop a policy for taking decisions to maximize the reward signals within a given environment. The environment is usually formulated as a Markov decision process. The state that the agent observed at time step $t$ is usually denoted as $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. The agent’s action at time step $t$ is denoted as $a_{t}$, which is determined by its policy $\pi:\mathcal{S}\rightarrow\mathcal{P(A)}$, where $\mathcal{P(A)}$ represents the set of probability distributions over the action space $\mathcal{A}$. In this framework, the policy is probabilistic, meaning that for any state $s\in\mathcal{S}$, $\pi(s)$ specifies a distribution over possible actions. The action taken ($a_{t}$) will let the agent move onto the next state ($s_{t+1}$). Such a transition is also mostly based on a probability distribution, denoted as $p(s_{t+1}|s_{t},a_{t})$. Accompanying the transmission from $s_{t}$ to $s_{t+1}$, the agent receives a reward signal, denoted as $r_{t}$, which is determined from a reward function ($r$). We can further denote the reward received at $t$ as such, $r_{t}=r(s_{t},a_{t},s_{t+1})$. The agent learns the policy via trials and errors, aiming to maximize the cumulative discounted future reward, $\eta(\pi)$, based on its policy ($\pi$). With the reward discount factor ($\gamma\in(0,1]$), the initial state ($s_{0}$), and the episode length $T$, we can formulate $\eta(\pi)$ as such: $\eta(\pi)=\mathbb{E}_{\pi}\Bigl[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\,\Bigl|\,s_{0}\Bigr]$. Deep RL is an advancement in RL that incorporates deep neural networks, to handle high-dimensional input spaces more effectively. Among all the deep RL algorithms, NAF is trained via Proximal Policy Optimizations (PPO) (Schulman et al., 2017), a common policy gradient-based algorithm. Our NAF+ is built upon Volpp et al. (2019), which also applied PPO as the training algorithm.
@@ 3. Neural Acquisition Function (NAF)
NAF keeps GP as the surrogate model of the task, allowing it to model the properties of the current optimization based on the real-time observations. However, instead of hand-crafted heuristic-based acquisition functions, NAF employs a neural network to generate the acquisition values (as introduced earlier). In the original paper, NAF (the model itself) at timestep $t$ as $\alpha_{t,\theta}$, where $\theta$ indicates that it is parameterized by a neural network. Similar to standard acquisition functions, NAF takes the information of the GP model as the input. However, a GP is a continuous function, but NAF is a deep neural network, whose input has a fixed dimension. Following Volpp et al. (2019), the design space $\mathcal{X}$ is discretized into a set of points $x_{s}$}. For each candidate $x$, NAF receives the GP’s predicted mean $\mu_{t}(x)$ and standard deviation $\sigma_{t}(x)$ as input features. Since the GP evolves with incoming observations, these features are dynamically updated at every iteration. In the original implementation, the set of $x_{s}$ was selected dynamically at each step. Note, to simplify training and ensure stable input representation, in our NAF+, we adopt a fixed set of $x_{s}$ throughout the optimization process. Additionally, to help the model account for optimization progress and urgency, NAF includes a budget-aware input: the percentage of optimization completed, calculated as the current iteration index divided by the total allowed iterations. This allows NAF to modulate its exploration–exploitation behavior over time. The output of the neural network is a categorical distribution over the discretized candidates, where each probability corresponds to the predicted acquisition value of selecting a specific $x_{s}$. During training, the next design candidate is sampled from this distribution, and the selected candidate is evaluated in the next iteration. During deployment, NAF directly selects the design candidate that has the highest acquisition value (value generated from NAF). We refer the readers to the original paper for more details about NAF (Volpp et al., 2019).
@@ 4. Summary and Limitations of NAF
NAF is a deep-learning-based meta-BO method with a key advantage: it can be trained with unlimited data without incurring computational overhead at deployment. Unlike GP-based approaches whose complexity grows with each new observation, NAF remains efficient regardless of dataset size. This makes NAF particularly well-suited for our HOMI, where synthetic users can be generated in virtually unlimited numbers via model sampling. Through such large-scale training, NAF learns acquisition policies that generalize across a wide range of synthetic user characteristics.
However, NAF has critical limitations when applied to real-world, human-involved optimization tasks. First, NAF is inherently designed to optimize for a single, fixed objective. In practice, many interactive systems involve multi-objective trade-offs, such as balancing speed and accuracy in input tasks, or comfort and visibility in interface layout. NAF lacks the flexibility to adjust its behavior when design objectives or user priorities shift. Second, NAF, as a deep learning model, performs well only when test-time tasks fall within the distribution of training tasks. NAF is likely to fail if real-world user behavior deviates significantly from the synthetic training distribution (i.e., constitutes an out-of-distribution task), which is likely given that human behaviors are inherently noisy and diverse.
@@ 4.2. Overview of NAF+
To address the aforementioned challenges, we introduce NAF+, a new approach that builds upon NAF but incorporates two critical enhancements: the ability to dynamically adjust to multi-objective formulations, and fallback mechanisms for handling out-of-distribution users. These enhancements make NAF+ more flexible, generalizable, and better-suited for real-world human-in-the-loop applications. We assume that the target application involves multiple design objectives, which are eventually aggregated into a single scalar objective via a weighted-sum formulation; this is an approach commonly found in HCI work (Liao et al., 2024a; Bailly et al., 2013). Our contribution lies in training NAF+ such that it can support dynamic setting of objective weights during deployment without requiring any retraining, which allows the optimizer to respond to varying user needs or preferences on the fly. We also assume that while novel users (whose behavior significantly deviates from synthetic users) may appear during deployment, they can still benefit from standard BO mechanisms (e.g., exploration via typical acquisition functions like Expected Improvement). Therefore, NAF+ integrates a fallback mechanism to maintain robustness when facing novel, unseen users.
NAF+ has four key components to jointly achieve these enhancements: (1) GP-based surrogate model (GP): This component models the objective function based on the user’s observed performance data. It provides the predicted mean and variance necessary for computing acquisition values and remains the foundation for all downstream decisions. (2) Neural acquisition function (naf): This deep neural network replaces hand-crafted acquisition functions by learning a task-specific policy. Unlike the original NAF, our model is explicitly trained to condition on objective weights, enabling it to adjust its search strategy according to the prioritized trade-offs during deployment. To distinguish it from the original method, we refer to our version in lowercase: naf. Two more elements are introduced to mitigate the novel user challenge: (3) Expected Improvement (EI): A classical, well-established acquisition function used as a fallback. When the neural model performs poorly (e.g., due to encountering a novel user), EI provides a stable and exploration-oriented alternative. The acquisition values produced by naf and EI are combined via a weighted sum. (4) Novelty detector: This module estimates how novel the current user is with respect to the training distribution. Based on this assessment, it dynamically adjusts the weight between the neural acquisition score (naf) and the EI score. As novelty increases, more weight is placed on EI to ensure robust performance. In the below subsections, we will expand the details of each components.
@@ 4.3. Element 1: GP-based Surrogate model (GP)
The first component is the surrogate model of the target objective, reflecting the ongoing user. Following the standard BO framework, we employed a GP model to fit all the observed designs ($x$) and resulting objective function values ($y$). The information of GP (specifically, the predicted mean and variance at any given $x$) can be leveraged to compute acquisition values.

### v2

@@ 3. Human-in-the-Loop Optimization with Model-Informed Priors
Our first contribution is the introduction of a novel conceptual framework, which we term HOMI, and the corresponding steps of pretraining meta-BO with synthetic users. While NAF+ represents one possible realization, the HOMI framework is intended to be general and extensible, enabling future work to explore alternative implementations. Below, we outline the core components of the framework.
@@ 3.1. Core Elements of HOMI:
HOMI is an interactive framework built upon three key components: meta-Bayesian optimization (meta-BO), user models, and end-users. Within the HOMI framework, Meta-BO refers to a computational optimizer trained across a distribution of synthetic users operating on the same interaction. During the optimizer training phase, meta-BO interacts with synthetic users generated from selected, relevant user models. Since these interactions are simulated, the optimizer can perform an unlimited number of evaluations to learn generalized adaptation strategies. User models are parameterized functions (e.g., Fitts’ law) that simulate user behavior and performance. By sampling different model parameter settings, we can generate a diverse population of synthetic users, each representing different motor abilities, interaction speeds, or usage contexts. End-users are real human participants interacting with the system during the deployment phase. During this phase, meta-BO observes users’ real-time feedback, such as performance metrics, interaction preferences, or other behavioral data, and uses it to rapidly infer user-specific characteristics and identify optimal designs with as few evaluation iterations as possible.
@@ 3.2. General Steps in HOMI:
Figure 1 illustrates the four main steps in the HOMI framework. Note that the first three steps comprise the offline meta-BO training phase, while only the fourth step occurs during real-time deployment.
Step 1: Model selection. HOMI framework requires synthetic users with varying performances to train the optimizer. These synthetic users are generated by parameterized user models (Purificato et al., 2024), where varying model parameters correspond to different user abilities, devices, or usage contexts. For example, for typing and target selection, Fitt’s law is the established model for predicting the movement time, which can be parametrized to predict performance with different devices or different abilities. The key novelty of HOMI is leveraging these models to simulate a large range of typical user behaviors by systematically varying parameters across a range of values.
For many interaction tasks, such parameter ranges are provided in the literature. For pointing and target selection, prior work reports Fitts’ law parameters across devices and conditions (Soukoreff and MacKenzie, 2004; MacKenzie, 1992a). For touch input, studies provide Gaussian touch-error models with empirically validated parameters (Bi and Zhai, 2013; Orphanides and Nam, 2017). For text entry, typing models offer parameterized keystroke latencies and error distributions (Soukoreff and MacKenzie, 2003; MacKenzie and Soukoreff, 2002). Similarly, models for continuous motion, such as the Steering Law, include validated parameter sets (Accot and Zhai, 1997). Together, these models allow researchers to construct diverse synthetic users by sampling from known parameter ranges or making informed parameter choices, without collecting new user data or refitting models from scratch. In summary, the first step of HOMI is to identify task-relevant models that capture how design variations affect user performance.
Step 2: Synthetic user generation: Given selected models and parameter ranges, synthetic users are generated by sampling model parameters from the corresponding distributions. Each synthetic user represents a distinct set of user characteristics, leading to a diverse simulated population.
Step 3: Meta-BO training with synthetic interactions: Before deploying to the end-users, meta-BO interacts with these synthetic users to learn how to efficiently identify optimal designs across different user profiles. Since synthetic evaluations are inexpensive compared to human evaluations, this phase allows for large-scale training. The diversity of synthetic users further ensures the optimizer does not overfit to specific traits, but rather learns transferable strategies for personalization.
@@ 4.1. Preliminaries
1. Acquisition functions and surrogate models in BO
BO has two core components: a surrogate model and an acquisition function. The surrogate model fits the observed design points ($x$) and corresponding objective values ($y$), enabling prediction of the posterior mean and variance of the objective function at any candidate $x$. The acquisition function utilizes the surrogate model’s predictions to compute “acquisition values” across the design space, and the candidate with the highest value is selected for evaluation. EI is one of the most widely adopted acquisition functions in BO due to its simplicity, closed-form expression, and robust performance in balancing exploration and exploitation. Formally, EI at a candidate point $x$ is defined as:
(1) $\text{EI}(x)=\mathbb{E}[\max(0,f(x)-f(x^{+}))],$
where $f$ is the objective function being optimized and $f(x^{+})$ denotes the best observed objective value so far.
A shared limitation of conventional acquisition functions is that they are not task-specific: they are fixed and do not learn from prior experience, even when tasks share similar characteristics. This missed opportunity limits their efficiency, especially in human-in-the-loop settings where evaluations are costly and rapid adaptation is required. This motivates a shift toward learned acquisition functions, in which the acquisition function is learned from prior data. The data is collected from similar users or, in our case, synthetic users, allowing it to generalize across users or tasks.
@@ 2. Deep reinforcement learning
Neural Acquisition Function (NAF) is an approach that allows the optimizer (meta-BO) to learn a task-specific acquisition function using a neural network. Instead of relying on a fixed acquisition function like Expected Improvement (EI), NAF aims to learn a search strategy that is tailored to a type of optimization problems which shared similar characteristics. A key challenge in learning such an acquisition function is the absence of ground truth labels. That is, there is no “correct” acquisition value to supervise the training. The predictions are inherently uncertain, noisy, and task-dependent. To overcome this, NAF is trained using reinforcement learning (RL). The neural network generates acquisition function values across the design space. The design candidate with the highest score is selected for evaluation. The outcome of this evaluation (i.e., the observed objective function value) is then converted into a reward signal: promising evaluation results yield positive rewards, while poor results generate negative feedback. Over time, the model learns to associate high acquisition values with promising candidates from its current status, refining its search policy. Viewed from an RL perspective, meta-BO acts as an RL agent, and the neural acquisition function (NAF) serves as its policy model, determining which action (i.e., which design candidate) the agent should take in each iteration given the GP’s status. This formulation enables task-level learning, where the acquisition policy improves through repeated interaction with synthetic training tasks.
@@
Here, we briefly define the key terms of RL to better introduce NAF. For a more detailed introduction to reinforcement learning, we refer readers to Sutton and Barto (2018) and Li (2017). RL is a machine learning paradigm that trains an agent to develop a policy for making decisions to maximize the reward signals within a given environment. The state that the agent observed at time step $t$ is usually denoted as $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. The agent’s action at time step $t$ is denoted as $a_{t}$, which is determined by its policy $\pi$. The action taken ($a_{t}$) will let the agent move onto the next state ($s_{t+1}$). Such a transition is also mostly based on a probability distribution, denoted as $p(s_{t+1}|s_{t},a_{t})$. Accompanying the transmission from $s_{t}$ to $s_{t+1}$, the agent receives a reward signal, denoted as $r_{t}$, which is determined from a reward function ($r$). We can further denote the reward received at $t$ as such, $r_{t}=r(s_{t},a_{t},s_{t+1})$. The agent learns the policy via trial and error, aiming to maximize the cumulative discounted future reward based on its policy.
Our NAF+ follows Volpp et al. (2019) to train the neural acquisition function using Proximal Policy Optimizations (PPO) (Schulman et al., 2017). We frame each optimization task as an RL episode: the GP surrogate defines the state, the NAF outputs acquisition scores as the policy, and selecting a design corresponds to taking an action. The observed objective value provides the reward. Across many synthetic tasks, PPO updates the NAF to learn an acquisition policy that generalizes beyond any specific function. Unlike prior work, our formulation incorporates human-centered requirements directly into the reward structure and training tasks, enabling the acquisition function to adapt to HCI scenarios. More details will be expanded in subsection 4.4 and Table 1. The results showed that NAF+ enables faster convergence than other methods, particularly showing statistically better performance than TAF in the second and third iterations.
@@ 3. Neural Acquisition Function (NAF)
NAF keeps GP as the surrogate model of the task, allowing it to model the properties of the current optimization based on the real-time observations. NAF employs a neural network to generate the acquisition values. In the original paper, NAF (the model itself) at timestep $t$ is $\alpha_{t,\theta}$, where $\theta$ indicates that it is parameterized by a neural network. Similar to standard acquisition functions, NAF takes the information of the GP model as the input. A GP defines a continuous predictive function over the entire design space and can be queried at arbitrarily many points. In contrast, our NAF is a deep neural network that requires a fixed-dimensional input vector to summarize the whole GP’s information. Following Volpp et al. (2019), the design space $\mathcal{X}$ is discretized into a set of points $x_{s}$}. For each candidate $x$, NAF receives the GP’s predicted mean $\mu_{t}(x)$ and standard deviation $\sigma_{t}(x)$ as input features. Since the GP evolves with incoming observations, these features are dynamically updated at every iteration. In the original implementation, the set of $x_{s}$ was selected dynamically at each step. To simplify training and ensure stable input representation, we adopt a fixed set of $x_{s}$ throughout the optimization process. Additionally, to help the model account for optimization progress and urgency, NAF includes a budget-aware input: the percentage of optimization completed, calculated as the current iteration index divided by the total allowed iterations. This allows NAF to modulate its exploration–exploitation behavior over time. The output of the neural network is a categorical distribution over the discretized candidates, where each probability corresponds to the predicted acquisition value of selecting a specific $x_{s}$. During training, the next design candidate is sampled from this distribution, and the selected candidate is evaluated in the next iteration. During deployment, NAF directly selects the design candidate that has the highest acquisition value (value generated from NAF). We refer the readers to the original paper for more details about NAF (Volpp et al., 2019).
@@ 4. Summary and Limitations of NAF
NAF has a key advantage: it can be trained with unlimited data without incurring computational overhead at deployment. This makes NAF particularly well-suited for our HOMI, where synthetic users can be generated in virtually unlimited numbers via model sampling. However, NAF has critical limitations when applied to real-world optimization tasks. First, NAF is inherently designed to optimize for a single, fixed objective. In practice, many interactive systems involve multi-objective trade-offs, such as balancing speed and accuracy in input tasks, or comfort and visibility in interface layout. NAF lacks the flexibility to adjust its optimization strategies when design objectives or user priorities shift. Second, NAF, as a deep learning model, performs well only when test-time tasks fall within the distribution of training tasks. When applied to HILO, that is, NAF can only perform well when the real user’s performance characteristics are highly aligned with the synthetic users’ performance profile used during training. NAF is likely to fail if real-world user performance deviates significantly from the synthetic users’ performances (i.e., constitutes an out-of-distribution task), which is likely given that human behaviors are inherently noisy and diverse.
@@ 4.2. Overview of NAF+
To address the aforementioned challenges, we introduce NAF+, a new approach that builds upon NAF but incorporates two critical enhancements: the ability to dynamically adjust to multi-objective formulations, and fallback mechanisms for handling out-of-distribution users. These enhancements make NAF+ more flexible, generalizable, and better-suited for real-world human-in-the-loop applications. We assume that the target application involves multiple design objectives, which are eventually aggregated into a single scalar objective via a weighted-sum formulation; this is an approach commonly found in HCI work (Liao et al., 2024a; Bailly et al., 2013). Our contribution lies in training NAF+ such that it can support dynamic setting of objective weights during deployment without requiring any retraining, which allows the optimizer to respond to varying user needs or preferences on the fly. We also assume that while novel users (whose behavior significantly deviates from synthetic users) may appear during deployment, they can still benefit from standard BO mechanisms (e.g., exploration via typical acquisition functions). Therefore, NAF+ integrates a fallback mechanism to maintain robustness when facing novel, unseen users.
As illustrated in Figure 2, NAF+ has four key components to jointly achieve these enhancements: (1) GP-based surrogate model (GP): This component models the objective function based on the user’s observed performance data. (2) Neural acquisition function (naf): This deep neural network replaces hand-crafted acquisition functions by learning a task-specific policy. Unlike the original NAF, our model is explicitly trained to condition on objective weights, enabling it to adjust its search strategy according to the prioritized trade-offs during deployment. To distinguish it from the original method, we refer to our version in lowercase: naf. Two more elements are introduced to mitigate the novel user challenge: (3) Expected Improvement (EI): An acquisition function used as a fallback. When the neural model performs poorly (e.g., due to encountering a novel user), EI provides a stable and exploration-oriented alternative. The acquisition values produced by naf and EI are combined via a weighted sum. (4) Novelty detector: This module estimates how novel the current user is with respect to the training distribution. Based on this assessment, it dynamically adjusts the weight between the neural acquisition score (naf) and the EI score. As novelty increases, more weight is placed on EI to ensure robust performance.
Note that two of the four elements (the GP and the EI) follow standard BO, and the naf component is adopted from prior work (Volpp et al., 2019). Our contribution lies in integrating these components and adapting them to the unique challenges of human-centered optimization, including dynamically weighted objectives and out-of-distribution users. We expand the details of each component below.
@@ 4.3. Element 1: GP-based Surrogate model (GP)
The surrogate model reflects the ongoing user’s performance. We follow the typical BO framework and employ a standard GP. Please refer to subsection 4.1 for a more detailed introduction.